OpenStudy (anonymous):

What is the vertex for the graph of y = -3x2 + 18x - 13? Answer (-3, -40) (-3, -94) (3, 23) (3, 14)

OpenStudy (mertsj):

The last one

OpenStudy (nenadmatematika):

y'=-6x+18=0 so x=3 plug it back into your equation and you'll get y=14 so your vetex is (3,14)

OpenStudy (anonymous):

I loved that nenad - I will teach my son :)

OpenStudy (nenadmatematika):

haha :D how old is he?

OpenStudy (anonymous):

17 - will be taking calc next year. He is in algebra 2 now , your method teaches to stones at once. :)

OpenStudy (nenadmatematika):

awesome! :D say hello to him :D

OpenStudy (anonymous):

@Maria, practice makes perfect! Vertex ( h, k) h = - (b/2a) = - ( 18/-6) = 3 f (3 ) = -3 (3)^2 + 18 (3) - 13 = ...